What is the direction of the acceleration of the block when it is at the point 0.35 m,y 0.65 m?
Answer #1

Solution,

On Differentiating wrt to x and y,

Fx = 2 x 5.75 x

Fy = - 3 x 3.75y^2

At point (0.35, 0.65)

Acceleration, ax = 2 x 5.75 x 0.35/0.04 = 100.63 m/s^2

Acceleration, ay = - 3 x 3.75 x 0.65^2/0.04 = - 118.83 m/s^2

Magnitude = sqrt(ax^2 + ay^2) = 155.71 m/s^2

Direction, theta = arctan (ay/ax) = 49.74 degree

So, theta = 360 - 49.74 = 310.26 degree

Comment in case any doubt please rate my answer ....